OpenStudy (anonymous):

PERCENT OF INCREASE? 6b to 9b

OpenStudy (anonymous):

50% increase

OpenStudy (anonymous):

9-6=3 then divide by 6 = .5(100) = 50%

OpenStudy (anonymous):

\[\frac{9b-6b}{6b}=\frac{3b}{6b}=\frac{1}{2}=0.5=50%\]

OpenStudy (anonymous):

50%